### v1

# Algorithm for finding local integrals of motion in quantum lattice models in the thermodynamic limit

Abstract
Local integrals of motion (LIOMs) play a key role in understanding the stationary states of closed macroscopic systems. They were found for selected integrable systems via complex analytical calculations. The existence of LIOMs and their structure can also be studied via numerical methods, which, however, involve exact diagonalization of Hamiltonians, posing a bottleneck for such studies. We show that finding LIOMs in translationally invariant lattice models or unitary quantum circuits can be reduced to a problem for which one may numerically find an exact solution also in the thermodynamic limit. We develop and implement a simple algorithm and demonstrate the efficiency of this method by calculating LIOMs and the Mazur bounds for infinite integrable spin chains and unitary circuits. Finally, we demonstrate that this approach correctly identifies approximate LIOMs in nearly integrable spin ladders.
Introduction – Closed macroscopic systems during time-evolution reach a stationary state that is determined by local conservation laws or, equivalently, by local integrals of motion (LIOMs). Generic systems have only a few conserved quantities (like Hamiltonian, particle number, magnetization) and evolve to a thermal Gibbs state [1, 2, 3]. Interacting integrable systems have extensive number of LIOMs and the relevant stationary states have the form that is consistent with the generalized Gibbs ensemble [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Identification of LIOMs is also the starting point for construction of the hydrodynamics or the generalized hydrodynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] in case of integrable models. Moreover, anomalous transport (or its anticipated absence) in various quantum systems has recently been linked to the presence of LIOMs. Most prominent examples concern strongly disordered systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], tilted chains with emergent conservation of the dipole moment [44] and systems exhibiting the Hilbert space fragmentation [45, 46, 44].
While the importance of LIOMs for the long-time dynamics appears unquestionable, in many cases their construction remains a challenging problem. Even for the most studied integrable interacting model, i.e. the XXZ chain, the complete set of LIOMs (that includes also quasilocal LIOMs) has been established analytically only recently [47, 48, 49, 50, 13, 51, 52]. In many cases, LIOMs are identified within complex calculations that are specific for the given studied model [53, 54, 55, 56, 57, 58, 59, 60]. Although there are model-independent approaches which allow to find LIOMs in integrable models [61, 62] or approximate LIOMs in nearly integrable systems [63], they are based on exact diagonalization of Hamiltonians. The latter poses serious limitations on the studied models and raises questions concerning the reliability of the finite-size scaling.
In this Letter, we show that finding LIOMs in translationally-invariant lattice models belongs to the same class of problems as the high-temperature expansion [64, 65, 66]. Namely, one can numerically find exact LIOMs and the corresponding Mazur bounds [67, 68] in the thermodynamic limit without diagonalizing the Hamiltonian. This holds true for quantum systems where the dynamics is set either by a Hamiltonian or by a unitary transformation, as is the case in unitary circuits. The geometry of the system is not important, and the only essential assumption concerns the dynamics that needs to be set by local transformations.
Method – We consider a system containing $L$ sites, which are enumerated by the index $l=1,\ldots,L$, and study extensive Hermitian operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, where the upper index, $s$, distinguishes between various operators. The densities, $\hat{O}^{s}_{l}$, are local and act on up to $M$ sites close to $l$. We introduce the Hilbert-Schmidt (HS) product $\langle\hat{O}^{s}\hat{O}^{p}\rangle={\rm Tr}(\hat{O}^{s\dagger}\hat{O}^{p})/(ZL)$ and the HS norm of operators, $||\hat{O}^{s}||^{2}=\langle\hat{O}^{s}\hat{O}^{s}\rangle$, where $Z$ is the dimension of the Hilbert space. We recall that the HS products of local operators, $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle$, can be exactly calculated for infinite lattice systems using the technique employed in the high-temperature expansion.
Our main assumption is that the dynamics is set either by a local Hamiltonian, $\hat{H}$, or by a repeated sequence of local unitary transformations $\hat{U}$. In the former case, we take $\hat{H}=\sum_{l}\hat{H}_{l}$, where $\hat{H}_{l}$ acts on $M_{H}$ sites. We look for local operators, $\hat{Q}$, which commute with the Hamiltonian, $[\hat{Q},\hat{H}]=0$ or, equivalently [69], we require vanishing of the norm of a Hermitian operator $\hat{Q}^{\prime}=i[\hat{Q},\hat{H}]$, i.e., $||\hat{Q}^{\prime}||^{2}=\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle=0$. We note that $\hat{Q}^{\prime}$ is still local, $\hat{Q}^{\prime}=\sum_{l}\hat{Q}^{\prime}_{l}$, where $\hat{Q}^{\prime}_{l}$ acts on up to $M+M_{H}-1$ sites. For this reason, $\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle$ can also be calculated exactly in the thermodynamic limit. In case of circuits built out of repeated sequence of local unitary transformations, the same reasoning can be applied for $\hat{Q}^{\prime}=\hat{U}^{\dagger}\hat{Q}\hat{U}-\hat{Q}$. I.e., $\hat{Q}^{\prime}$ is a local operator and vanishing norm of $\hat{Q}^{\prime}$ implies that $\hat{Q}$ is conserved. Both cases can be comprehended within a unified formulation. Namely, we introduce a function $f(\hat{O}^{s})$ which has the following properties: (i) $f$ is linear, (ii) vanishing of $||f(\hat{O}^{s})||$ is equivalent to conservation of $\hat{O}^{s}$; (iii) $f(\hat{O}^{s})$ is a local and Hermitian operator so that one may exactly calculate $\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$. If the dynamics is set by a Hamiltonian then one may choose $f(\hat{O}^{s})=i[\hat{O}^{s},\hat{H}]$, whereas $f(\hat{O}^{s})=\hat{U}^{\dagger}\hat{O}^{s}\hat{U}-\hat{O}^{s}$ can be used for unitary circuits. The specific choice of $f$ is irrelevant as long as it is consistent with conditions (i)-(iii).
As a next step, we consider a set containing $N_{O}$ orthonormal operators, $\langle\hat{O}^{s}\hat{O}^{p}\rangle=\delta_{sp}$. We aim to find all LIOMs which can be expressed as linear combinations of these operators. To this end, we build the matrix containing HS products $F_{sp}=\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$ and solve the eigenproblem,
$\sum_{sp}V_{\alpha s}F_{sp}\;(V^{T})_{p\beta}=\lambda_{\alpha}\delta_{\alpha\beta},$ (1)
where $V$ is an orthogonal matrix and $F$ is a matrix built in the basis (of size $N_{O}$) containing all $\hat{O}^{s}$. Since $f$ is a linear function, one may introduce rotated operators
$\hat{A}^{\alpha}=\sum_{s}V_{\alpha s}\hat{O}^{s},$ (2)
and rewrite Eq. (1)
$\langle f(\hat{A}^{\alpha})f(\hat{A}^{\beta})\rangle=\lambda_{\alpha}\delta_{\alpha\beta},$ (3)
from which one finds the norm $||f(\hat{A}^{\alpha})||=\sqrt{\lambda_{\alpha}}$. The rotated operators are local and orthogonal, $\langle\hat{A}^{\alpha}\hat{A}^{\beta}\rangle=\delta_{\alpha\beta}$. Consequently, all $\hat{A}^{\alpha}$ corresponding to the eigenvalue $\lambda_{\alpha}=0$ represent orthogonal LIOMs, identified here via the norm $||f(\hat{A}^{\alpha})||=0$. To distinguish these LIOMs from other operators introduced in Eq. (2), we denote LIOMs as $\hat{Q}^{\alpha}$ where $\alpha=1,\dots,N_{L}$ and $N_{L}$ is the number of LIOMs given by the degeneracy of the eigenvalue $\lambda_{\alpha}=0$ in Eq. (1). We keep the notation $\hat{A}^{\alpha}$, for the remaining operators, $\alpha=N_{L}+1,...,N_{O}$ with $\lambda_{\alpha}>0$.
As a final step, we show that the solution of Eq. (1) identifies all LIOMs that can be represented via linear combinations of $\hat{O}^{s}$. To this end, we consider an operator $\hat{B}=\sum_{s}b_{s}\hat{O}^{s}$, $||\hat{B}||>0$, that is orthogonal to all $\hat{Q}^{\alpha}$ obtained from Eq.(1). We show that $||f(\hat{B})||>0$ implying that $\hat{B}$ is not conserved. Using the transformation from Eq. (2), such operator can be expressed in the rotated basis
$\hat{B}$ $=$ $\sum_{\alpha=1}^{N_{L}}\left(\sum_{s}b_{s}V_{\alpha s}\right)\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\left(\sum_{s}b_{s}V_{\alpha s}\right)\hat{A}^{\alpha}$ (4)
$=$ $\sum_{\alpha=1}^{N_{L}}\tilde{b}_{\alpha}\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{b}_{\alpha}\hat{A}^{\alpha}.$
The first term in Eq. (4) vanishes since we assume that $\hat{B}$ is orthogonal to all $\hat{Q}^{\alpha}$. Then, one finds the squared norm
$||f(\hat{B})||^{2}$ $=$ $\sum_{\alpha,\beta=N_{L}+1}^{N_{O}}\tilde{b}_{\alpha}\tilde{b}_{\beta}\langle f(\hat{A}^{\alpha})f(\hat{A}^{\beta})\rangle$ (5)
$=$ $\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{b}^{2}_{\alpha}\lambda_{\alpha}>0,$
where we have used Eq. (3). The inequality in Eq. (5) follows from that all $\lambda_{\alpha}$ are positive for $\alpha>N_{L}$ and some $\tilde{b}_{\alpha}$ are nonzero (otherwise $||\hat{B}||=0$, contradicting one of the assumptions).
Solving the eigenproblem in Eq. (1) gives not only the LIOMs but also the Mazur bounds for all operators $\hat{O}^{s}$ used in the construction of the $F$ matrix. Using Eq.(2) one finds that the infinite-time ($t$) correlation functions are bounded by the matrix $V$,
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\rm d}t\,\langle\hat{O}^{s}(t)\hat{O}^{s}\rangle\geq\sum_{\alpha=1}^{N_{L}}\frac{\langle\hat{O}^{s}\hat{Q}^{\alpha}\rangle^{2}}{\langle\hat{Q}^{\alpha}\hat{Q}^{\alpha}\rangle}=\sum_{\alpha=1}^{N_{L}}V^{2}_{\alpha s}\,.$ (6)
The basis of operators in spin-1/2 systems – The first step is to choose the set of operators, $\hat{O}_{s}$, for which we then construct the matrix in Eq. (1). Any specific knowledge concerning LIOMs may help reduce the set of operators that are used for its construction. Without such knowledge, one may implement a brute-force approach and systematically include all operators with growing support. In such case, one may use at each site, $l$, the set containing four operators: the identity matrix, $\hat{\sigma}^{0}_{l}$, and the Pauli matrices $\hat{\sigma}^{x}_{l}$, $\hat{\sigma}^{y}_{l}$, $\hat{\sigma}^{z}_{l}$. These operators are orthogonal, $\langle\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}\rangle=\delta_{s,s^{\prime}}/L$, so they may be used for constructing extensive operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, supported on $m$ sites
$\hat{O}^{s}_{l}=\hat{O}^{\left(s_{1},...,s_{m}\right)}_{l}=\hat{\sigma}^{s_{1}}_{l+1}\hat{\sigma}^{s_{2}}_{l+2}...\hat{\sigma}^{s_{m}}_{l+m}.$ (7)
We assume that the first, $\hat{\sigma}^{s_{1}}_{l+1}$, and the last, $\hat{\sigma}^{s_{m}}_{l+m}$, spin operators are other than $\hat{\sigma}^{0}$. All operators with maximal support $M$ (i.e., generated for all $m\leq M$) are orthogonal $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle=\delta_{l,l^{\prime}}\delta_{s,s^{\prime}}/L$. For fixed $M$, one has in total $N_{O}=3\times 4^{M-1}$ extensive operators which are orthonormal $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$. We note that all translationally invariant operators supported on up to $M$ sites can be expressed as linear combinations of $\hat{O}^{s}$ thus one may easily calculate the HS product of such operators. Finally, we recall that product of two $\hat{\sigma}^{s}_{l}$-operators is proportional to other such operator, $\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}=\alpha_{s^{\prime\prime}}\hat{\sigma}^{s^{\prime\prime}}_{l}$, where $\alpha_{s^{\prime\prime}}\in\{1,\pm i\}$. This enables a straightforward calculation of $f(\hat{O}^{s})$, which we implement using the software package from Ref. [70].
XXZ chain – The integrable XXZ chain appears as an obvious system for testing our method. The brute-force approach, as described above, allows to account for all operators supported on up to $M=8$ sites yielding $N_{O}=3\times 4^{7}\sim 5\times 10^{4}$ basis operators $\hat{O}^{s}$. To go beyond this limitation, we resolve the symmetries of the Hamiltonian (time-reversal, parity, conservation of total $\hat{S}^{z}$), and note that LIOMs can be found independently in each symmetry-sector. Then, it is convenient to use $\{\hat{\sigma}^{0}_{l},\hat{\sigma}^{+}_{l},\hat{\sigma}^{-}_{l},\hat{\sigma}^{z}_{l}\}$ on-site basis, where $\hat{\sigma}^{\pm}_{l}=(\hat{\sigma}^{x}_{l}\pm i\hat{\sigma}^{y}_{l})$. This basis is orthogonal but non-Hermitian so we first build the operators as in Eq. (7) and then introduce the Hermitian counterparts, $\hat{O}^{{s}}_{l,R}=\hat{O}^{{s}}_{l}+\hat{O}^{{s}\dagger}_{l}$ and $\hat{O}^{{s}}_{l,I}=i(\hat{O}^{{s}}_{l}-\hat{O}^{{s}\dagger}_{l})$. We numerically solve the eigenproblem in Eq. (1) for $f[\hat{O}^{s}_{R(I)}]=i[\hat{O}^{s}_{R(I)},\hat{H}_{\text{XXZ}}]$ with
$\hat{H}_{\text{XXZ}}=\sum_{l=1}^{L}\frac{1}{2}\left[\hat{\sigma}_{l}^{+}\hat{\sigma}_{l+1}^{-}+\hat{\sigma}_{l}^{-}\hat{\sigma}_{l+1}^{+}\right]+\Delta\hat{\sigma}_{l}^{z}\hat{\sigma}_{l+1}^{z}.$ (8)
Fig. 1(a) shows the lowest $\lambda_{\alpha}$ obtained for $\Delta=3/4$ in the basis build from Pauli matrices (the Pauli strings basis). One observes that the eigenvalue $\lambda_{\alpha}=0$ is $M$-fold degenerate. The corresponding eigenvectors $\hat{Q}^{\alpha}$ represent the well-known integrals of motion of the XXZ chain. In particular $\hat{Q}^{1}$ obtained for $M=1$ is the total $S^{z}$. Additional $\hat{Q}^{2}$ for $M=2$ is the Hamiltonian and subsequent $\hat{Q}^{3}$ for $M=3$ is the energy current, see Appendix A in the End Matter for more details. Fig. 1(b) shows analogous results obtained in the symmetry-resolved basis, $\{\hat{O}^{{s}}_{R}\}$, containing operators that conserve total $\hat{S}^{z}$ and are even under time-reversal as well as the spin-flip symmetries. It reveals $M/2$ LIOMs in this symmetry sector, and we reach support up to $M=10$. In Ref. [71], we provide a simple code, based on Ref. [70], that generates these LIOMs. In the End Matter (Appendix B) we also demonstrate that investigating the dependence of $\lambda_{\alpha}$ and $V_{\alpha s}$ on $M$ allows one to identify quasilocal LIOMs [47, 48] as well as their contributions to the Mazur bound (e.g. for the spin current operator).
Figure 1: Lowest eigenvalues obtained from Eq. (1) for the XXZ model with $\Delta=3/4$ and different operator bases. (a) Pauli strings basis yields exactly $M$ LIOMs with support $m\leq M$. (b) Symmetry resolved basis, restricted to operators commuting with total $\hat{S}^{z}$ and even under time-reversal and parity transformations. One obtains $M/2$ LIOMs in this symmetry sector.
XXX unitary circuit – As a second example, we consider the simplest integrable unitary quantum circuit, obtained from the integrable trotterization of XXX Heisenberg chain [72]. It is a brickwall quantum circuit, constructed from a two-site unitary matrix
$\hat{U}_{l,l+1}(\delta)=\frac{1+i\frac{\delta}{2}(1+\hat{h}_{l,l+1})}{1+i\delta}\,,$ (9)
with $\hat{h}_{l,l+1}=1/2\left[\hat{\sigma}^{+}_{l}\hat{\sigma}^{-}_{l+1}+\hat{\sigma}^{-}_{l}\hat{\sigma}^{+}_{l+1}\right]+\hat{\sigma}^{z}_{l}\hat{\sigma}^{z}_{l+1}$ being the local density of the XXX Hamiltonian. The one-step unitary propagator is then defined as
$\hat{\mathcal{U}}(\delta)=\prod_{l=1}^{L/2}\hat{U}_{2l-1,2l}(\delta)\;\;\prod_{l=1}^{L/2}\hat{U}_{2l,2l+1}(\delta)$ (10)
It consists of two half-steps, acting first on all even-odd, and next on all odd-even nearest neighbors. Although $\hat{\mathcal{U}}(\delta)$ alone is manifestly nonlocal, $\hat{\mathcal{U}}(\delta)^{\dagger}\hat{O}^{s}\hat{\mathcal{U}}(\delta)$ is still a local operator.
Figure 2: Solution of Eq. (1) for the XXX quantum circuit defined via elementary gate in Eq. (9) with $\delta=0.5$. (a) shows the eigenvalues and (b) shows the structure of LIOMs visualized via the Mazur bound from Eq. (6), $B_{s}=\sum_{\alpha=1}^{N_{L}}V^{2}_{\alpha s}$.
For every LIOM $\hat{Q}_{n}$ from continuous dynamics, we obtain two LIOMs $\hat{Q}_{n}^{\pm}(\delta)$ in discrete dynamics, related to each other by single-site translation [19]. Both $\hat{Q}_{n}^{\pm}(\delta)$ are supported on $12n+$ 1 sites, and in the limit $\delta\to 0$ reproduce XXX LIOMs. This observation allows us to choose an appropriate operator basis for our algorithm. We start with the same local basis as in the XXZ chain case, but consider extensive operators invariant under two-site translations, $\hat{O}^{{s},-}=\sum_{l}\hat{O}^{{s}}_{2l-1}$ and $\hat{O}^{{s},+}=\sum_{l}\hat{O}^{{s}}_{2l}$. Due to the staggered structure of the propagator $\hat{\mathcal{U}}(\delta)$, we have to relax the assumption that the first operator in the support is different from $\hat{\sigma}^{0}_{l}$. Here, we avoid trivial LIOMs related to magnetization conservation by removing $m=1$ operators from the basis.
Fig. 2(a) shows eigenvalues obtained from Eq. (1). Since LIOMs correspond to the degenerate subspace with $\lambda_{\alpha}=0$, they can be arbitrarily rotated. One may fix this rotation to restore the form of LIOMs known from the literature or according to any other criterion. However the Mazur bound from Eq. (6), $B_{s}=\sum_{\alpha=1}^{N_{L}}V^{2}_{\alpha s}$, is independent of such rotations. Therefore, in Fig. 2(b) we show $B_{s}$ for various operators $\hat{O}^{s}$ which are grouped according to their support $m$. We see two LIOMs for $M=3$. Extending the maximal support from $M=3$ to $M=4$ neither yields additional LIOMs nor alters (up to numerical precision $\sim 10^{-12}$) the structure of LIOMs obtained for $M=3$. The next two LIOMs emerge for $M=5$ in agreement with Ref. [72].
Nearly integrable systems– Finally, we study a two-leg spin ladder
$\hat{H}_{\text{ladder}}$ $=\sum_{y=1}^{2}\hat{H}_{y}+U\sum_{l=1}^{L}\hat{\sigma}^{z}_{l,y=1}\hat{\sigma}^{z}_{l,y=2}\;,$ (11)
$\hat{H}_{y}$ $=\sum_{l=1}^{L}\frac{1}{2}\left[\hat{\sigma}^{+}_{l,y}\hat{\sigma}^{-}_{l+1,y}+\hat{\sigma}^{-}_{l,y}\hat{\sigma}^{+}_{l+1,y}\right]+\Delta\hat{\sigma}^{z}_{l,y}\hat{\sigma}^{z}_{l,y}\;,$
where $\hat{\sigma}^{s}_{l,y}$ is spin operator at site $l$ in leg $y=1,2$. For each leg, we use operators that are even under spin-flip and conserve the total $\hat{S}^{z}$. Then, we consider all possible products of densities that are odd under time-reversal, $\hat{O}^{{s}}_{l,y=1,R}\hat{O}^{{s^{\prime}}}_{l,y=2,I}$ and $\hat{O}^{{s}}_{l,y=1,I}\hat{O}^{{s^{\prime}}}_{l,y=2,R}$, and build the basis of extensive operators, $\hat{O}^{ss^{\prime}}$. This symmetry sector includes energy currents of the uncoupled XXZ chains.
We have chosen the Hamiltonian (11) because it is integrable for either $\Delta=0$ (Bose-Hubbard chain) or $U=0$ (two XXZ chains). It has been shown in Ref. [73] that the nearly integrable regime in this model is peculiar in that the relaxation times may be extremely long. However, the model turned out to be too challenging to perform the finite-size scaling and to estimate the relaxation times in the thermodynamic limit. Our present approach does not provide any quantitative information about the time-dependence. However, one may expect that operators, $\hat{A}^{\alpha}$, with the longest relaxation times correspond to the smallest norms $||i[\hat{H},\hat{A}^{\alpha}]||^{2}=\lambda_{\alpha}$, see also Ref. [69]. Upon adding a perturbation, $g\hat{H}^{\prime}$, to the integrable Hamiltonian, $\hat{H}$, one finds that $||i[\hat{H}+g\hat{H}^{\prime},\hat{Q}^{\alpha}]||^{2}=g^{2}||i[\hat{\hat{H}}^{\prime},\hat{Q}^{\alpha}]||^{2}\propto g^{2}$. Such $g$-dependence is generally expected for the relaxation rates. Therefore, we test whether $\lambda_{\alpha}$ can be considered as an estimate of the relaxation rate of $\hat{Q}^{\alpha}$.
Fig. 3(a) shows the smallest eigenvalue, $\lambda_{1}$, obtained from Eq. (1) taking $M=3$ per each leg. One observes that $\lambda_{1}$, vanishes on both integrable lines, $U=0$ and $\Delta=0$, while for small $U$ and $\Delta$ its value is determined by the product $\Delta U$ (see the isolines). The corresponding operator, $\hat{A}^{1}$, has dominant contributions from the symmetric combination of the XXZ energy currents in both legs, $\hat{Q}^{+}=\hat{Q}^{3}_{y=1}+\hat{Q}^{3}_{y=2}$, and the three-site Hubbard current-like LIOM, $\hat{Q}^{3}_{H}$. For small $U$ and $\Delta$, $\hat{Q}^{+}$ is almost the same as $\hat{Q}^{3}_{H}$. Fig. 3(b) shows that $\lambda_{1}$ accurately follows the dependence $\lambda_{1}\propto(U\Delta)^{2}$. Perturbative arguments and the Boltzmann equations [73] suggest the same dependence for the relaxation rate, $\Gamma^{+}$, of the operator $\hat{Q}^{+}$. Most importantly, $\lambda_{1}$ agrees even quantitatively with $\Gamma^{+}$ estimated in Ref.  [73] from the Lanczos propagation method for a ladder with $L=14$ rungs. In the End Matter (Appendix C), we discuss the results for other slow mode that corresponds to $\lambda_{2}$.
Figure 3: (a) Heatmap of the smallest eigenvalue obtained from Eq.(1) for spin ladder defined in Eq.(11). Continuous white curves represent isolines. (b) Cross-section of results from (a) along the red line. $\Gamma^{+}$ shows the relaxation rate of $\hat{Q}^{+}$ estimated in Ref. [73] from Lanczos propagation method.
Concluding remarks – We have derived and implemented a numerical approach that provides exact translationally-invariant LIOMs and the corresponding Mazur bounds for infinite quantum systems for which the dynamics is set by either local Hamiltonian or by local unitary transformations. We are able to obtain the exact LIOMs because our approach, similarly to Refs. [74, 69], is targeting the short-time dynamics (as encoded, e.g., in the commutators $[\hat{O},\hat{H}]$) under which local operator $\hat{O}$ remains local. It is an important advantage over approaches focused on the infinite-time properties that are encoded in the diagonal matrix elements of observables calculated in the basis of the eigenstates of the Hamiltonian. In the latter case, the finite-size effects seem unavoidable.
Like all numerical methods, our approach also has limitations. This bottleneck concerns the number of local observables, $N_{O}\sim 10^{4}-10^{5}$, from which LIOMs are built by solving the eigenproblem in Eq. (1). In case of spin systems, we still considered all operators supported on up to 10 consecutive sites. On the one hand, one may expect that finding very complex LIOMs may be problematic. On the other hand, such LIOMs only have a limited influence on the dynamics of physically relevant observables, which are typically represented by few-body local operators.
Our approach is not based on diagonalization of the Hamiltonian or a unitary matrix. Therefore, it can be applied to problems that are hardly accessible to other numerical methods, e.g., when one studies problems beyond one-dimensional chains. In particular, it also identifies quasilocal LIOMs as well as LIOMs in the Hubbard model. Furthermore, we have demonstrated that this approach also applies to nearly integrable systems, where it singles out slowly decaying observables and provides reasonable estimates of the corresponding relaxation rates. It is rather obvious that one may also carry out the same calculations for systems that are not translationally invariant. In the latter case, although one cannot study infinite systems anymore, the accessible system sizes are still much larger than is the case within approaches that are based on the diagonalization procedure.
Acknowledgments. J.P. acknowledges support by the National Science Centre (NCN), Poland via project 2023/49/N/ST3/01033. M.M. acknowledges support by the National Science Centre (NCN), Poland via project 2020/37/B/ST3/00020.